OpenStudy (anonymous):

Simplify: cos^2(x) - 1 + sin^2(x) + sin(x) I have rewritten it to be: cos^2(x)+ sin^2(x) + sin(x) - 1 = 1-1+sin(x), so I get sinx as a result. Am I thinking right?

OpenStudy (anonymous):

@phi, just need a quick look. Thank you.

OpenStudy (phi):

yes, that looks good \[ \cos^2(x) - 1 + \sin^2(x) + \sin(x) \\ (1- \sin^2(x)) - 1 + \sin^2(x) + \sin(x) \\1- 1 - \sin^2(x) + \sin^2(x) + \sin(x) \\\sin(x) \]

OpenStudy (anonymous):

Thank you!